# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ V-A2 Utility of HuUs
where $U^{\mathrm{H}}_{\mathbbm{j},0}$ corresponds to local execution as a baseline.
V-A3 Utility of MaUs
An MaU $u^{\mathrm{On}}_{\mathbbm{j}}\in\bm{\mathcal{U}}^{\mathrm{A}}$ is associated with a latency deadline $D_{\mathbbm{j}}>0$. Its total edge-side latency at $s^{\mathrm{On}}_{\mathbbm{i}}$ is $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}$ in (13). To model their deadline-aware behavior, we adopt a soft-deadline valuation as $G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,\frac{\big[D_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}{D_{\mathbbm{j}}},$ where $x]_{+}=\max\{x,0\}$, and $v_{\mathbbm{j}}>0$ is the MaU’s task reward. Intuitively, the MaU only accrues reward when the deadline is met, while partial lateness leads to a proportional loss of reward. Similar to HuUs, the MaU’s energy saving from enjoying edge services is $c^{\mathrm{A},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$. Let $\mathbb{V}^{\mathrm{A}}_{e}$ refer to MaU’s valuation coefficient on energy saving. MaU’s utility is given by
$U^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}+\mathbb{V}^{\mathrm{A}}_{e}\,c^{\mathrm{A},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}-p_{\mathbbm{j},\mathbbm{i}},\qquad U^{\mathrm{A}}_{\mathbbm{j},0}=0.$ (15)
@@ V-A3 Utility of MaUs
Note that both HuU and MaU utilities in (14) and (15) share the same congestion component $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$ in (12). This common structure is pivotal for constructing an ordinal potential function that aligns individual incentives with system-wide performance in the next subsection.
V-A4 Cost of SP
@@
For SP $s^{\mathrm{On}}_{\mathbbm{i}}$, the aggregate computing load induced by its associated users is $\mathbbm{x}_{\mathbbm{i}}=\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}r_{\mathbbm{v}}$. We model the SP-side computation/operation cost as $K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}})=\omega_{1}\,e^{\mathrm{comp}}_{\mathbbm{i}}\;\frac{\mathbbm{x}_{\mathbbm{i}}}{f^{\mathrm{On}}_{\mathbbm{i}}}+\omega_{2}c^{\mathrm{hard}}$, where $e^{\mathrm{comp}}_{\mathbbm{i}}$ is the average computing power consumption of $s^{\mathrm{On}}_{\mathbbm{i}}$. Although a SP may collect service payments $\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}p_{\mathbbm{v},\mathbbm{i}}$, these payments are pure monetary transfers among agents and therefore cancel out when aggregating the system-level welfare. Consequently, they do not appear in the potential function constructed later and only affect how the welfare is divided among participants rather than the overall efficiency.
@@ V-B Ordinal Potential Function
We construct an ordinal potential[20, 13] for the online game under the soft-deadline utilities of MaUs, leveraging the unified congestion structure established above. The key observation is that all users (HuUs and MaUs) experience congestion through the same factorized computing delay as (12). Then, for MaU $u^{\mathrm{On}}_{\mathbbm{j}}$ assigned to $s^{\mathrm{On}}_{\mathbbm{i}}$, the soft-deadline valuation is
$G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,\frac{\big[D_{\mathbbm{j}}-\big(t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}\big)\big]_{+}}{D_{\mathbbm{j}}},\qquad[x]_{+}=\max\{x,0\}.$ (16)
Since $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}$ in (12) is increasing in $y_{\mathbbm{i}}$, heavier congestion directly reduces $G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}$. In other words, load accumulation at a SP gradually erodes the MaU’s reward for meeting its deadline.
For HuUs, the utility in (14) can be decomposed into a congestion-independent part and a congestion-dependent penalty, given by
$U^{\mathrm{H}}_{\mathbbm{j},\mathbbm{i}}=\underbrace{\mathbb{V}^{\mathrm{H}}_{t}\!\big(t^{\mathrm{loc}}_{\mathbbm{j}}-t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}\big)+\mathbb{V}^{\mathrm{H}}_{e}\!\big(E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}\big)}_{S_{\mathbbm{j}}(\mathbbm{i})\;\text{(independent of }y_{\mathbbm{i}}\text{)}}\;-\;\underbrace{\mathbb{V}^{\mathrm{H}}_{t}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}(y_{\mathbbm{i}})}_{\text{congestion penalty}}\;-\;p_{\mathbbm{j},\mathbbm{i}}.$ (17)
where $S_{\mathbbm{j}}(\mathbbm{i})$ collects all terms that do not depend on the congestion load $y_{\mathbbm{i}}$, while the second term captures how queueing-induced delay degrades HuU utility.
Under soft deadlines, MaU utilities admit the same structural form. By linearizing $G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}$ around typical delay levels (see Appx. D-D), the MaU utility can be rewritten as
$U^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=\bar{S}_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}(y_{\mathbbm{i}})\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (18)
where $\bar{S}_{\mathbbm{j}}(\mathbbm{i})$ collects the non-congestion terms (soft-deadline base reward and energy saving), and $\kappa_{\mathbbm{j}}>0$ reflects the MaU’s effective sensitivity to congestion. This linearized form is standard in congestion and potential-game analysis and preserves the ordinal structure of preferences, even if the original soft-deadline utility is mildly nonlinear.
Considering HuUs and MaUs, the utility of SD $u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}}$ taking action $\bm{X}^{\mathrm{On}}\in\{0,1\}$ can be unified as
@@
where $S_{\mathbbm{j}}(0)=0$, and where $\tau_{\mathbbm{i}}(\cdot)$ is relevant only for $\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}$ (local execution does not suffer edge-side congestion). This unified form makes explicit that all heterogeneity is captured by user-specific coefficients $(S_{\mathbbm{j}},\kappa_{\mathbbm{j}},\theta_{\mathbbm{j},\mathbbm{i}})$ on top of a shared congestion mechanism.
Let $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ denote the joint assignment profile, where $\pi_{\mathbbm{j}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ and $\pi_{\mathbbm{j}}=0$ stands for local execution, while $\pi_{\mathbbm{j}}=\mathbbm{i}$ means offloading to $s^{\mathrm{On}}_{\mathbbm{i}}$. Given $\bm{\pi}$, the effective congestion load and the physical computing load at $s^{\mathrm{On}}_{\mathbbm{i}}$ are
@@
which drive, respectively, the delay inflation via $\tau_{\mathbbm{i}}(\cdot)$ and the SP-side cost via $K_{\mathbbm{i}}(\cdot)$.
Thus, we define the following potential function:
$\Phi(\bm{\pi})=\underbrace{\sum_{\mathbbm{j}:\,\pi_{\mathbbm{j}}\neq 0}S_{\mathbbm{j}}(\pi_{\mathbbm{j}})}_{\text{user-side non-congestion benefits}}$ (21)
@@
$-\underbrace{\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}\int_{0}^{\,y_{\mathbbm{i}}(\bm{\pi})}\tau_{\mathbbm{i}}(z)\,dz}_{\text{integrated time-congestion cost}}-\underbrace{\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}K_{\mathbbm{i}}\!\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)}_{\text{SP computation/operation cost}}.$
where, the first term aggregates all non-congestion benefits obtained by users that offload to the edge. The second term integrates the congestion function $\tau_{\mathbbm{i}}(\cdot)$ from zero to the realized load $y_{\mathbbm{i}}(\bm{\pi})$, and thus acts as a “convex penalty” for overloading SPs. Under the linear response $\tau_{\mathbbm{i}}(y)=y/f^{\mathrm{On}}_{\mathbbm{i}}$, this integral simplifies to $y_{\mathbbm{i}}^{2}(\bm{\pi})/\big(2f^{\mathrm{On}}_{\mathbbm{i}}\big)$, which is strictly convex in $y_{\mathbbm{i}}$. The last term accounts for SP-side computing and operational costs, ensuring that the potential internalizes both SD-side and SP-side effects.
Consider a unilateral deviation where exactly one SD switches from SP $s^{\mathrm{On}}_{\mathbbm{i}}$ to SP $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$ (or to local execution). The change in $\Phi(\bm{\pi})$ has three components:
  • The increment in the first term of (21) equals the change in that SD’s non-congestion benefit, $S_{\mathbbm{j}}(\cdot)$.
  • The increment in the second term is determined by the difference between the integrated congestion before and after the deviation. Since $\tau_{\mathbbm{i}}(\cdot)$ is increasing, this difference has the same sign as the change in that SD’s individual congestion penalty, $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$, up to a scalar factor; this is a standard “integral–pointwise sandwich” argument in congestion games.
  • The increment in the third term reflects SP-side cost variation due to the convexity of $K_{\mathbbm{i}}(\cdot)$ w.r.t. the physical load $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$.
Therefore, whenever a SD changes its action in a way that strictly increases its own utility (disregarding pure payment transfers), the induced change in $\Phi(\bm{\pi})$ has the same sign. In this sense, $\Phi(\bm{\pi})$ serves as an ordinal potential under the soft-deadline setting [^5] Payment terms $p_{\mathbbm{j},\mathbbm{i}}$ are pure monetary transfers between SPs and SDs, and are excluded from $\Phi(\bm{\pi})$; they affect the distribution of SW but not the total SW. A short justification is provided in Appx. D-I..
V-C Properties of the Ordinal Potential
We now formalize $\Phi(\bm{\pi})$ in (21) as an ordinal potential for the online game under soft-deadline utilities. The proof follows a standard single-user deviation argument and relies on three ingredients: (i) separability of non-congestion terms $S_{\mathbbm{j}}(\cdot)$, (ii) a sandwich bound connecting the pointwise congestion penalty in user utilities with the integrated congestion cost in $\Phi(\bm{\pi})$, and (iii) monotonicity/convexity of the SP-side cost $K_{\mathbbm{i}}(\cdot)$. Full details are moved to Appendices D-CD-G.
Theorem 1 (Ordinal potential).
Consider utilities in (19) with soft-deadline MaUs and the potential $\Phi(\bm{\pi})$ in (21). For any SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and any unilateral deviation $\pi_{\mathbbm{j}}=\mathbbm{i}\to\mathbbm{i}^{\prime}$, if $U_{\mathbbm{j},\mathbbm{i}^{\prime}}-U_{\mathbbm{j},\mathbbm{i}}>0$, then $\Phi\big(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}^{\prime}\big)-\Phi\big(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}\big)>0$. Hence, $\Phi(\bm{\pi})$ is an ordinal potential for the considered game[13].
Since each SD possesses a finite action set, any sequence of unilateral best-response updates yields a strictly increasing trajectory of the potential function $\Phi(\bm{\pi})$ and therefore converges in finitely many iterations to a pure-strategy NE. Consequently, every local maximizer of $\Phi(\bm{\pi})$ corresponds to a pure-strategy NE. A formal proof is provided in Appx. D-H.
V-D Problem Formulation

### v2

@@ V-A2 Utility of HUs
where the value of $U^{\mathrm{H}}_{\mathbbm{j},0}$ corresponds to local task execution.
V-A3 Utility of MUs
While HUs mainly care about the joint latency–energy tradeoff, MUs (e.g., robots or sensing devices) are typically driven by task deadlines and reliability requirements. We therefore adopt a slightly different utility structure for MUs. Specifically, an MU $u^{\mathrm{On}}_{\mathbbm{j}}\in\bm{\mathcal{U}}^{\mathrm{M}}$ is associated with a latency deadline $\tau^{\mathrm{ddl}}_{\mathbbm{j}}$, while its total edge-side latency at $s^{\mathrm{On}}_{\mathbbm{i}}$ is $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}$ in (12). To capture the fact that the MU primarily values meeting this deadline, we adopt a soft-deadline valuation
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,{\big[\tau^{\mathrm{ddl}}_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}\times({\tau^{\mathrm{ddl}}_{\mathbbm{j}}})^{-1},$ (14)
where $x]_{+}=\max\{x,0\}$ and $v_{\mathbbm{j}}>0$ denotes the MU’s task reward. Intuitively, the MU obtains the full reward $v_{\mathbbm{j}}$ if the deadline is met, whereas excessive latency leads to a loss of reward (once the deadline is violated, the reward becomes zero). In particular, “soft-deadline” means that this reward decreases smoothly as latency approaches the deadline, instead of dropping abruptly from full reward to zero at a hard threshold.
In addition to timeliness, an MU also benefits from saving local computing energy by task offloading. Similar to HUs, the MU’s energy saving when utilizing edge services is $c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$, where $E^{\mathrm{loc}}_{\mathbbm{j}}$ is the local execution energy and $E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ is the uplink transmission energy. Let $\mathbb{V}^{\mathrm{M}}_{e}$ denote the MU’s valuation coefficient on energy saving. Combining deadline satisfaction, energy saving, and the service payment $p_{\mathbbm{j},\mathbbm{i}}$, the MU’s utility when offloading its tasks to $s^{\mathrm{On}}_{\mathbbm{i}}$ is given by:
$U^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}+\mathbb{V}^{\mathrm{M}}_{e}\,c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}-p_{\mathbbm{j},\mathbbm{i}},\qquad U^{\mathrm{M}}_{\mathbbm{j},0}=0,$ (15)
where $U^{\mathrm{M}}_{\mathbbm{j},0}$ corresponds to local task execution.
@@ V-A3 Utility of MUs
Note that both HU and MU utilities in (13) and (15) inherently share the same congestion component $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$ in (11). This shared structure will be crucial for aligning individual incentives with system-wide performance via an ordinal potential function in our subsequent analysis (Sec. V-B).
V-A4 Cost of SPs
Having specified how HUs and MUs (i.e., SDs) evaluate edge services, we now focus on the SPs’ perspective. As more SDs offload their tasks to an SP, the SP consumes more energy and experiences higher operating costs, which must be reflected in its payoff and, ultimately, in the system-level SW. Subsequently, for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$, given its aggregated computing load $\mathbbm{x}_{\mathbbm{i}}=\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}r_{\mathbbm{v}}$, we model its computation/operation cost as follows:
$K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}})=\omega_{1}\,e^{\mathrm{comp}}_{\mathbbm{i}}{\mathbbm{x}_{\mathbbm{i}}}\big/{f^{\mathrm{On}}_{\mathbbm{i}}}+\omega_{2}c^{\mathrm{hard}},$ (16)
@@ V-A4 Cost of SPs
where $e^{\mathrm{comp}}_{\mathbbm{i}}$ is the average computing power consumption of $s^{\mathrm{On}}_{\mathbbm{i}}$, and $c^{\mathrm{hard}}$ captures hardware- and maintenance-related costs ($\omega_{1},\omega_{2}\geq 0$ are weighting coefficients). In (16), the term proportional to $\mathbbm{x}_{\mathbbm{i}}/f^{\mathrm{On}}_{\mathbbm{i}}$ reflects that processing a larger workload leads to higher energy and operational expenditure. It is worth noting that although an SP may collect service payments $\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}p_{\mathbbm{v},\mathbbm{i}}$, these payments are monetary transfers among SP and SDs and therefore cancel out when aggregating the system-level SW. Consequently, they do not appear in the potential function constructed later in Sec. V-B and only affect how the SW is divided among SPs and SDs.
In a nutshell, in our considered human–machine ecosystem, SPs resources are limited and SDs (i.e., HUs/MUs) act selfishly, where each SD tries to choose the execution mode of its tasks (local vs. offloading) and SP that maximizes its own utility. Directly coordinating all the coupled decisions across SPs and SDs via a centralized SW maximization formulation is computationally prohibitive and poorly scalable. This motivates our following game-theoretic treatment in which we capture the system performance by a carefully crafted potential function that is aligned with individual incentives and allows for decentralized handling of resource allocations across SPs and SDs.
@@ V-B Ordinal Potential Function
In order to analyze our online resource scheduling game and to enable simple distributed best-response dynamics with provable convergence, we first construct an ordinal potential $\Phi(\bm{\pi})$ that aggregates SD-side benefits and SP-side costs. This construction leverages the unified congestion structure in (11) together with the above-described deadline-aware MUs’ utilities.
For an MU $u^{\mathrm{On}}_{\mathbbm{j}}$ served by $s^{\mathrm{On}}_{\mathbbm{i}}$, the time-related part of its payoff was modeled by the soft-deadline valuation in (14) as $G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,{\big[\tau^{\mathrm{ddl}}_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}\times{(\tau^{\mathrm{ddl}}_{\mathbbm{j}})^{-1}}$. This valuation, together with energy saving and the payment term, determines the MU’s overall utility in the online game defined in (15). We note that since the computing delay $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}$ in (11), which is inherently encoded in (15), is an increasing function of the congestion load $y_{\mathbbm{i}}$, heavier congestion directly reduces $G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}$ and the utility in (11). This monotone relationship between congestion and MU valuation allows us to align individual best responses with the system-wide potential in the subsequent analysis.
For HUs, the utility in (13) can be decomposed into a congestion-independent valuation part and a congestion-dependent penalty, as follows:
[…]
where $S_{\mathbbm{j}}(\mathbbm{i})$ collects all valuation terms that do not depend on the congestion load $y_{\mathbbm{i}}$, while the second term captures how queueing-induced delay degrades the HU’s utility. Note that the payment $p_{\mathbbm{j},\mathbbm{i}}$ is also independent of $y_{\mathbbm{i}}$, but is written separately because it represents a monetary transfer, which will cancel out at the system level in our subsequent potential analysis.
MU utilities also admit the same structural form: by linearizing $G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}$ around typical delay levels (see Appx. E-D), the MU utility can be rewritten as:
$U^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=\bar{S}_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}(y_{\mathbbm{i}})\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (18)
where $\bar{S}_{\mathbbm{j}}(\mathbbm{i})$ gathers the non-congestion valuation terms (soft-deadline valuation base reward and energy saving), and $\kappa_{\mathbbm{j}}>0$ reflects the MU’s effective sensitivity to congestion. Similar to (17), $p_{\mathbbm{j},\mathbbm{i}}$ is separated from $\bar{S}_{\mathbbm{j}}(\mathbbm{i})$ in (18) even though it does not depend on $y_{\mathbbm{i}}$, so that monetary transfers can be excluded from the system-level potential.
Considering both HUs and MUs, the utility of SD $u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}}$ when it is served by $s^{\mathrm{On}}_{\mathbbm{i}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ can be unified as follows:
@@
where $\mathbbm{i}=0$ corresponds to local task execution at SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and $S_{\mathbbm{j}}(0)=0$. In (19), $S_{\mathbbm{j}}(\mathbbm{i})$ encodes the non-congestion valuation terms (latency/energy saving or soft-deadline valuation), $\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}$ captures how sensitive $u^{\mathrm{On}}_{\mathbbm{j}}$ is to the congestion at $s^{\mathrm{On}}_{\mathbbm{i}}$, and $\tau_{\mathbbm{i}}(\cdot)$ is the server-side congestion function that only applies when $\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}$ (local execution does not impact edge-side congestion). The term $p_{\mathbbm{j},\mathbbm{i}}$ represents the payment charged to $u^{\mathrm{On}}_{\mathbbm{j}}$ and, while independent of congestion, is written explicitly as it will later be treated as a monetary transfer in the system-level analysis. The unified representation in (19) captures SD (i.e., HU/MU) heterogeneity through $(S_{\mathbbm{j}},\kappa_{\mathbbm{j}},\theta_{\mathbbm{j},\mathbbm{i}})$, whereas all SDs are coupled through the same congestion mechanism $\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$ at each SP.
To describe how individual decisions interact at the system-level, we collect all SD assignments into the joint assignment profile [^4] Given $\bm{\pi}$, the binary indicators $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}$ are induced by $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=\mathbbm{1}_{\{\pi_{\mathbbm{j}}=\mathbbm{i}\}},\quad\mathbbm{i}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ ($\mathbbm{1}_{\{\cdot\}}$ denotes the binary indicator function), so that the vector $\bm{X}^{\mathrm{On}}_{\mathbbm{j}}$ and the scalar action $\pi_{\mathbbm{j}}$ are equivalent representations of $u^{\mathrm{On}}_{\mathbbm{j}}$’s decision. $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$, where $\pi_{\mathbbm{j}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ denotes the choice of $u^{\mathrm{On}}_{\mathbbm{j}}$: $\pi_{\mathbbm{j}}=0$ means that the task is processed locally at $u^{\mathrm{On}}_{\mathbbm{j}}$, while $\pi_{\mathbbm{j}}=\mathbbm{i}$ indicates task offloading to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. Given a profile $\bm{\pi}$, we can then compute, for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$, the effective congestion load and the physical computing load as follows:
@@
where $y_{\mathbbm{i}}(\bm{\pi})$ feeds into the congestion-induced delay via $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}(\bm{\pi})\big)$ in (11), and $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ determines the SP-side computation/operation cost through $K_{\mathbbm{i}}\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)$ in (16). Thus, any unilateral change of $\pi_{\mathbbm{j}}$ only modifies the loads of the involved SPs in (20), and these updated loads propagate to SD delays and SP costs through (11) and (16), respectively.
In order to analyze the resulting game and to show that simple myopic best-response updates can converge, we now construct an ordinal potential function that aggregates both SD-side benefits and SP-side costs as follows:
@@
$\hskip-5.69054pt{\footnotesize\begin{aligned} \Phi(\bm{\pi}){=}\hskip-19.91692pt&\underbrace{\sum_{\mathbbm{j}:\,\pi_{\mathbbm{j}}\neq 0}S_{\mathbbm{j}}(\pi_{\mathbbm{j}})}_{\text{user-side non-congestion benefits}}\hskip-8.53581pt-\underbrace{\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}\int_{0}^{\,y_{\mathbbm{i}}(\bm{\pi})}\hskip-4.2679pt\tau_{\mathbbm{i}}(z)\,dz}_{\text{aggregated congestion cost}}-\hskip-8.53581pt\underbrace{\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}K_{\mathbbm{i}}\!\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)}_{\text{SP computation/operation cost}}\hskip-5.69054pt,\end{aligned}}\hskip-14.22636pt\vskip 2.84526pt$ (21)
where the first term sums up the non-congestion valuation of all SDs that choose to offload, the second term integrates the congestion function $\tau_{\mathbbm{i}}(\cdot)$ from zero to the realized load $y_{\mathbbm{i}}(\bm{\pi})$ across SPs, and the last term captures the cost of running edge resources. By jointly accounting for SD-side benefits, congestion penalties, and SP-side costs, $\Phi(\bm{\pi})$ serves as a system-level measure whose variations mirror individual utility changes; this coupling is key to proving that our online resource scheduling game is an ordinal potential game and that distributed better/best-response dynamics converge to a pure-strategy NE. In particular, consider a unilateral deviation where one SD $u^{\mathrm{On}}_{\mathbbm{j}}$ switches from SP $s^{\mathrm{On}}_{\mathbbm{i}}$ to another SP $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$ (or to local task execution). The resulting change in $\Phi(\bm{\pi})$ can be decomposed as follows:
  • User-side non-congestion term: The first term in (21) only depends on the non-congestion valuation $S_{\mathbbm{j}}(\cdot)$ of the deviating SD, while all other SDs keep the same assignment. Hence, its increment is exactly the change in that SD’s non-congestion benefit, i.e., $S_{\mathbbm{j}}(\pi_{\mathbbm{j}}^{\text{new}})-S_{\mathbbm{j}}(\pi_{\mathbbm{j}}^{\text{old}})$.
  • Congestion term: The second term aggregates the integrated congestion at each SP through $\int_{0}^{y_{\mathbbm{i}}(\bm{\pi})}\tau_{\mathbbm{i}}(z)\,dz$, where $\tau_{\mathbbm{i}}(y)$ is an increasing function of the congestion load $y$ (recall $\tau_{\mathbbm{i}}(y)=y/f^{\mathrm{On}}_{\mathbbm{i}}$). When $u^{\mathrm{On}}_{\mathbbm{j}}$ moves from $s^{\mathrm{On}}_{\mathbbm{i}}$ to $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$ (or from/to local execution), the loads $y_{\mathbbm{i}}$ and $y_{\mathbbm{i}^{\prime}}$ change by fixed amounts. Because $\tau_{\mathbbm{i}}(\cdot)$ and $\tau_{\mathbbm{i}^{\prime}}(\cdot)$ are increasing functions and $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}>0$, the sign of the change in the integrated congestion term matches the sign of the change in that SD’s congestion penalty $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$; this follows the standard integral–pointwise comparison used in congestion games[27].
  • SP-side cost term: The third term in (21) captures the SP-side computation/operation cost $K_{\mathbbm{i}}\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)$, where $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ is the aggregate computing load at $s^{\mathrm{On}}_{\mathbbm{i}}$. A deviation by $u^{\mathrm{On}}_{\mathbbm{j}}$ only changes the loads at $s^{\mathrm{On}}_{\mathbbm{i}}$ and $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$, and, due to the convex and increasing behavior of $K_{\mathbbm{i}}(\cdot)$ with respect to $\mathbbm{x}_{\mathbbm{i}}$, the sign of the change in this term is aligned with the way that the SD’s action changes the corresponding SP-side cost.
Putting the above components together, it can be immediately construed that whenever $u^{\mathrm{On}}_{\mathbbm{j}}$ changes its action in a way that strictly increases its own utility (excluding the payment term $p_{\mathbbm{j},\mathbbm{i}}$), the resulting change in $\Phi(\bm{\pi})$ has the same sign as its utility improvement. In other words, $\Phi(\bm{\pi})$ is an ordinal potential for our online resource scheduling game under the soft-deadline valuations: any unilateral utility-improving deviation by an SD strictly increases $\Phi(\bm{\pi})$. This property is crucial as it implies that simple distributed better-/best-response dynamics monotonically climb the potential and therefore converge in finitely many steps to a pure-strategy NE. In particular, we mathematically prove that $\Phi(\bm{\pi})$ in (21) is an ordinal potential, so that any unilateral utility-improving move strictly increases $\Phi(\bm{\pi})$ and any better-/best-response dynamics converge to a pure-strategy NE (see Appx. C for details); the complete technical development, including the matrix reformulation, the integral–pointwise “sandwich” inequality, and the finite-improvement property, is deferred to Appx E.
V-C Problem Formulation